OpenStudy (anonymous):

b^2 + 49 How do I tell if both terms are squares?

OpenStudy (anonymous):

b^2 is pretty clearly the square of b yes?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

how about 49?

OpenStudy (anonymous):

see iff the other part of the equation is the product of the same number, here it is 7

OpenStudy (anonymous):

and how did you find 7 as your answer. I see that it's correct, but how do I find it?

OpenStudy (asadkarim7):

keep practsing with the squared numbers uptil 256 and you would kinda memorize it.

OpenStudy (asadkarim7):

7---7^2 = 49

OpenStudy (anonymous):

Sry, still not quite getting a beginner's method here...

OpenStudy (anonymous):

if a the number can be divisible by 1 number and it have the same as the answer

OpenStudy (anonymous):

\[i^2=1\] \[2^2=4\] \[3^2=9\] \[4^2=16\] \[5^2=25\] \[6^2=36\] \[7^2=49\] \[8^2=64\] \[9^2=81\] question just asks you to recognize 49 as the square of 7